the area of a rectangle is 165ft^2. the length is 4ft greater than the width. find the dimensions of the rectangle.
Answer by checkley75(3666)

W*L=165, L=W+4
W(W+4)=165
W^2+4W-165=0
(W-11)(W+15)=0
W-11=0
W=11 ANSWER.
PROOF
11(11+4)=165
11*15=165
165=165